OpenStudy (anonymous):

equation for the given points (-2,8), (-6,0)?

OpenStudy (austin1):

the hittest

OpenStudy (anonymous):

Point–slope form y - y_1 = m( x - x_1 ) equation is 2x-y+12=0

OpenStudy (anonymous):

Since the equation of a slope is a=y2-y1/x2-x1 you can then plug and chug the two given points 0-8/(-6)-(-2) ( The answer from that is 2).therefor a =2 The y-intercept is found by a*x1+b=y1 (once again plug & chug) 2*(-2)+b=12. In conclusion the equation that would best fit these two given points would be y=2x+12